# VideoGLaMM [Uncaptioned image]: A Large Multimodal Model for Pixel-Level Visual Grounding in Videos

### v1

@@ 1 Introduction
  • We assess the performance of VideoGLaMM across diverse tasks spanning grounded conversation generation, visual grounding, and referring video segmentation, where it achieves state-of-the-art performance.
@@ 2 Related work
Visual Grounding. Recently Grounded LMMs [35, 38, 9, 19, 52, 47, 49] have made significant strides in enhancing visual and language comprehension and excel in complex localization tasks. These models demonstrate proficiency in tasks such as referring expression comprehension and image segmentation, highlighting the advanced image understanding capabilities of LLMs. Approaches such as [35, 9, 47] primarily focus on creating a language-based context for visual grounding. In contrast, [52] integrates visual elements with language, while [19] leverages vision-language embeddings to produce segmentation masks. Additionally, [38] is adept at generating natural language responses linked with object segmentation masks, facilitating detailed visual-textual interactions. However, these models are limited to image-based applications and do not extend to video understanding. Recently, Munasinghe et al. [31] incorporates audio transcripts alongside visual and textual data for a more detailed video understanding. However, it combines pre-trained modules that cannot be trained end-to-end. Although effective in video understanding, the lack of fine-grained spatial-temporal modeling in [31] hinders the precise understanding and location of segments in detail. To this end, we propose a novel fine-grained alignment strategy to align language instruction across both spatial and temporal dimensions. Our approach facilitates a more finegrained video understanding via precise visual grounding.
@@ 3.1 Overview
For end-to-end spatio-temporal alignment, we train VideoGLaMM on our proposed fine-grained benchmark dataset, consisting of 38k video-QA pairs with 83k objects and roughly 671k masks. During training, we finetune the LoRA parameters of the LLM, along with V→L and L→V adapters. This approach seamlessly combines spatial and temporal data through an improved alignment mechanism and a precise grounding framework, enhancing the model’s capability for visual grounding and understanding. The proposed model can perform a diverse range of tasks, including grounded conversation generation, visual grounding, and referring video segmentation.
@@ 3.2 Architecture
The overall architecture of our VideoGLaMM consists of the following components: (i) Spatio-Temporal Dual Encoder, (ii) Dual Alignment V-L Adapters for image and video features, (iii) Large Language Model (LLM), (iv) Promptable Pixel Decoder. Below we provide a detailed description and also explain how the input is handled in each of these components.
@@
$f_{g}=\mathcal{F}_{g}(V),V\in\mathbb{R}^{T\times H\times W\times C}$ (1)
Meanwhile, for extracting video features, we use segment-wise Sampling following [30] to obtain fine-grained temporal cues. Given an input video $V\in\mathbb{R}^{T\times H\times W\times C}$, we divide it into $K$ segments, where each segment consists of $s=\frac{T}{K}$. The video encoder $\mathcal{F}_{h}$, operates on low-resolution video segments $V_{k}\in\mathbb{R}^{s\times H\times W\times C}$ yielding global features that provide segment-wise temporal context.
$f_{h}=\mathcal{F}_{h}(V_{k}),V_{k}\in\mathbb{R}^{s\times H\times W\times C}$ (2)
@@
$Z_{g}$ $=\mathcal{W}_{g}(f_{g})$ (3)
$Z_{h}$ $=\mathcal{W}_{h}(f_{h})$
@@
Figure 3: Proposed Semi-automatic Annotation Pipeline. Our dataset for grounded conversation generation (GCG) is built from three video dataset types: i) Videos having masks only: Object patches are extracted from video frames using masks and processed by the Gemini model for initial object descriptions, which are then refined to produce detailed object captions. These refined captions and masks are used again with the Gemini model to create dense, grounded captions. ii) Videos having bbox annotations and captions: Frames are first processed with a Video-LMM to generate a comprehensive caption which is combined with the original caption and fed to GPT-4o to obtain dense grounded captions. Masks are generated using frames and ground-truth bounding boxes with the SAM model. iii) Videos having object bboxes and referring expressions: Frames, bounding boxes, and referring expressions are input to GPT-4o for dense grounded captions, while masks are generated by feeding frames and bounding boxes to the SAM model.
@@ 3.3 Training Strategy
We train VideoGLaMM end-to-end in a single stage. As stated above, we use a dual encoder consisting of separate image and video encoders for processing spatial and temporal inputs to obtain local and global features, respectively. These encoders are initialized with weights if strong encoders. During the training, we keep the encoders fixed and only train the V→L adapters $\mathcal{W}_{g}$ and $\mathcal{W}_{h}$ associated with these encoders. These adapters are used to project the spatio-temporal visual features in the space of LLM and align the two modules. The spatio-temporal encoder is kept frozen and only the V→L adapters are updated. The textual features from the last layer of LLM, rich in spatial and temporal cues, are projected into the space of the pixel decoder using a multi-layer projection L→V adapter $\mathcal{W}_{p}$. For the LLM, we keep its weights frozen and only finetune LoRA [14] parameters during training. The frame encoder associated with the pixel decoder is instantiated with pre-trained weights and produces multi-scale features from the input video frames. We keep the frame encoder and pixel decoder frozen and only train the L→V adapter layer. We optimize the output of the LLM by minimizing the cross entropy CE objective between the autoregressively obtained text output and dense grounded ground-truth caption. For the output of mask decoder, we optimize the intersection over union (IOU) between the predictions of mask decoder and ground-truth masks denoted as $\mathcal{L}_{masked}$. The total loss is the sum of CE loss and masked loss.
@@ 4 Our Benchmark & Annotation Pipeline
b) Videos with Bounding Box annotations and Captions: Fig. 3(b) shows the annotation process for the videos having both captions and object bounding box (Bbox) annotations. To obtain the corresponding dense grounded caption, the video frames are first passed to an open-source Video-LMM [26] to obtain a detailed caption, which is fed along with the reference ground truth caption to GPT-4o mini [33] to obtain the final dense grounded caption. The Bbox annotations are used as a prompts to a SAM model [18] which takes the video frames as input and provides the masks corresponding to the objects.
@@ 5 Experimental Setup
Implementation details. Our spatio-temporal dual encoders follow the design of image and video encoders from VideoGPT+ [30]. For the image encoder, we use a pretrained CLIP ViT-L/14 (336 × 336)[36] model, and for the temporal encoder, we select the pretrained encoder of InternVideov2 (224 × 224) [48]. The V→L projectors are initialized with the weights of MLP adapter from VideoGPT+ [30]. The LLM is instantiated with Phi3-Mini-3.8B [1] weights. Both the frame encoder and pixel decoder are initialized with SAM2 [40] encoder-decoder weights. The training (Sec. 3.3) is carried out end-to-end on 4 Nvidia A100 40GB GPUs with a distributed training based on DeepSpeed [39].
Datasets. We train the model on our proposed grounded conversation (GCG) dataset containing 38k grounded video-QA triplets along with 83k objects and 671k fine-grained masks. Our proposed benchmark is derived from YTVOS [16], BURST [3] ActivityNet entities [54], Refer-YTVOS [43], MEVIS [12], VidSTG [51] and HCSTVG [44] datasets. During training, we also include a variety of other image and video datasets with our proposed benchmark dataset for more robust alignment. Our choice of image datasets include: ADE20K [53], COCO-Stuff [8], LVIS-PACO [37], refCOCO, refCOCO+, refCLEF, refCOCOg [15], LLaVA-Instruct-150k [24], ReasonSeg [19] and GranDf [38]. For video-datasets we include train samples from: Refer-DAVIS17 [17], VideoInstruct100K [29].
Tasks. We evaluate VideoGLaMM on three challenging tasks: grounded conversation generation (GCG), visual grounding, and referring video segmentation, in zero-shot settings. For grounded conversation generation, we curate a separate dataset of 308 refined video-QA triplets containing 826 objects and 22,762 fine-grained masks, following our proposed annotation pipeline. For Visual Grounding, we evaluate our model on two challenging datasets: VidSTG [51] and HC-STVG [44]. VidSTG dataset consists of 5693 videos paired with multiform sentences (both interrogative and declarative). In the case of referring video segmentation, we leverage the MeViS [12] validation dataset consisting of 140 videos and 2,236 referring sentences. All the results on MeViS dataset are obtained via official CodaLab evaluation suite.
@@
5.1 Grounded Conversation Generation
The Grounded Conversation Generation (GCG) task aims to provide video-level detailed captions with specific phrases directly tied to corresponding segmentation masks in the video frames. For example, “<An adult> in white clothes holds a <cup> in the room", as shown in the first row of Fig. 4, features how each bracketed phrase is anchored to a unique segmentation mask. This creates a densely annotated caption that aligns textual descriptions with visual regions in the frames, enriching the video’s contextual interpretation. To obtain GCG output, we query the model with the following sample prompt: “Provide a detailed description of the image. Respond with interleaved segmentation masks for the corresponding parts of the answer.” The model generates a detailed caption along with interleaved segmentation masks, employing the format “<p>An adult woman in brown</p><SEG> is talking to another <p>adult man wearing jacket</p><SEG>” a shown in the third row of Fig. 4. We use special tokens, namely <p>, </p> and <SEG>, to delineate the start and end of each phrase and its corresponding region mask, respectively.
Our proposed GCG test set is meticulously constructed as explained in Sec. 4, capturing fine-grained annotations. The test set contains 308 refined video-QA triplets with 826 objects and 22,762 fine-grained masks. As shown in Table 1, our proposed Video-GLaMM performs better in generating detailed captions containing references to objects in the video frames, as is evident from high METEOR, CIDEr and CLAIR scores. Regarding the quality of masks, Video-GLaMM consistently outperforms baselines in terms of mIOU and Recall scores, signifying a higher overlap with ground-truth masks. Fig. 4 further shows the qualitative visualizations of VideoGLaMM on GCG samples.
@@ 5.2 Referring Video Segmentation
For referring video segmentation, the output should be grounded as per the given phrase, pointing towards specific instances in the video. Given a sentence or referring expression containing a specific object instance, the goal is to localize the object instances present across the video frames. This task operates in an open vocabulary setting, assessing the model’s ability to localize objects both spatially and temporally. Given a referring phrase expression Phrase, we prompt the model using the following instruction prompt to obtain the instance masks: “What is {Phrase} in this video? Respond with segmentation masks". As seen from Table 2, both the region Jaccard $\mathcal{J}$ and boundary F-measure $\mathcal{F}$ are high in the case of VideoGLaMM, significantly outperforming the baselines. Similarly, the mean score $\mathcal{J\&F}$ follows the same trend. Additionally, the scores corresponding to VideoLISA are reported with post-processing step. Notably, VideoLISA involves an additional post-processing step to boost performance. Therefore, we further fine-tune the VideoGLaMM on the task of referring segmentation post epoch 30 until epoch 40. Clearly, VideoGLaMM outperforms the Video-LISA (post-processed) on both $\mathcal{J}$ and $\mathcal{F}$, including the mean $\mathcal{J\&F}$.
@@
Table 2: Performance comparison of VideoGLaMM on referring expression segmentation: VideoGLaMM shows superior performance on motion grounding and segmenting referring objects in the videos. VideoGLaMM-FT, which is further finetuned on the task of referring segmentation, shows improvements over post-processed VideoLISA.
@@ 5.3 Visual Grounding
To quantitatively assess VideoGLaMM’s visual grounding capability, we conduct quantitative evaluations on the benchmark datasets derived from the test set of VidSTG dataset as stated in the Tasks section above. The visual grounding task measures the adeptness of the model at correlating textual descriptions with visual elements in the video, a critical aspect of contextual comprehension. This ability is crucial for applications that integrate continuous visual data with language. The output of this task is refined masks that correlate with the caption caption. To obtain the visual grounding output, we query the model with interrogative captions. For these captions, the prompt format follows “{caption} Please respond with a segmentation masks.". Table 3 shows VideoGLaMM’s improved visual grounding precision as it outperforms the baselines and demonstrates its fine-grained understanding.
@@ 5.4 Ablation studies
Effect of Spatio-Temporal Dual Encoder. We employ separate image and video encoders to process spatial and temporal information. While spatial processing induces local information, temporal processing includes a global view of video in the model. Both are necessary from the perspective of grounding. To verify the effectiveness of using the dual spatio-temporal encoder, we conduct an ablation study where we measure the effectiveness of each encoder for grounded conversation generation task (see Table 4). We notice that using only an image encoder gives suboptimal results, as we notice a drop in both the localization and captioning metrics. Using only video branch leads to the highest mIOU; however, relatively lower METEOR, CIDEr, and CLAIR scores. To get the best of both worlds, i.e., optimal mIOU and good conversational abilities, VideoGLaMM uses both image and video branches.
@@
Spatial vs Spatio-temporal Pixel decoder. Pixel decoder in VideoGLaMM can operate in two configurations. The first configuration processes video frames individually, ignoring temporal consistency. The second configuration employs both spatial and temporal branches for spatio-temporal context. Table 5 demonstrates the impact of adding a temporal branch alongside the spatial branch in our grounded conversation generation dataset. Results indicate that using only the spatial configuration reduces performance, with a nearly 3% drop in mIOU scores compared to the spatio-temporal configuration. Similarly, metrics like METEOR, CIDEr, and CLAIR also show a decline.
@@
Effect of number of frames for Pixel Decoder. The Pixel decoder receives the raw input frames as supervision signals through the frame encoder. However, Pixel decoder can only take one frame as input at a time, which doesn’t offer temporal information. To provide more temporal supervision, we instead feed the Pixel decoder multiple frames to enhance its temporal understanding. This allows it to learn semantic information that generalizes across frames. Table 6 shows the performance when 4 and 8 frames are input to the decoder. We observe that while the mIOU with 8 frames is slightly lower compared to 4 frames, the conversational quality measured by METEOR, CIDEr and CLAIR is higher. Hence, to achieve a decent mIOU with higher conversational output, we stick to 8 frames as input.
@@
6 Limitations and Future Work
Our GCG dataset plays a key role in enhancing the model’s grounding capabilities. While we validated annotations manually, some noise may still be present. Also, each scene contains several objects and the video descriptions do not exhaustively cover all objects in the scenes. A higher-quality densely annotated set could further boost model performance but would require substantial annotation resources. Additionally, VideoGLaMM struggles with objects of varying granularities, likely due to limited representation in the training data. Another improvement is to extend VideoGLaMM for longer videos, as the current GCG dataset mainly focuses on short-medium duration clips.

### v2

@@ 1 Introduction
  • We assess the performance of VideoGLaMM across diverse tasks spanning grounded conversation generation (GCG), visual grounding, and referring video segmentation, where it achieves state-of-the-art performance.
@@ 2 Related work
Visual Grounding. Recently Grounded LMMs [36, 39, 9, 20, 54, 49, 51] have made significant strides in enhancing visual and language comprehension and excel in complex localization tasks. These models demonstrate proficiency in tasks such as referring expression comprehension and image segmentation, highlighting the advanced image understanding capabilities of LLMs. Approaches such as [36, 9, 49] primarily focus on creating a language-based context for visual grounding. In contrast, [54] integrates visual elements with language, while [20] leverages vision-language embeddings to produce segmentation masks. Additionally, [39] is adept at generating natural language responses linked with object segmentation masks, facilitating detailed visual-textual interactions. However, these models are limited to image-based applications and do not extend to video understanding. Recently, [32] incorporates audio transcripts alongside visual and textual data for a more detailed video understanding. However, it combines pre-trained modules that cannot be trained end-to-end, which results in lack of fine-grained spatiotemporal modeling. Similarly, [5] introduced a new video grounding model, but their architecture employs only a spatial encoder-decoder setup and does not address the GCG task. To this end, we propose VideoGLaMM, which leverages a novel fine-grained alignment strategy to align language instruction across both spatial and temporal dimensions, facilitating more finegrained video understanding.
@@ 3.1 Overview
For end-to-end spatio-temporal alignment, we train VideoGLaMM on our proposed fine-grained benchmark dataset. During training, we finetune the LoRA parameters of the LLM, along with V→L and L→V adapters. This approach seamlessly combines spatial and temporal data through an improved alignment mechanism and a precise grounding framework, enhancing the model’s capability for visual grounding and understanding.
@@ 3.2 Architecture
The overall architecture of our VideoGLaMM consists of following components: (i) Spatio-Temporal Dual Encoder, (ii) Dual Alignment V-L Adapters, (iii) Large Language Model (LLM), (iv) Pixel Decoder. Below we provide a detailed description and working of each of component.
@@
$f_{g}=\mathcal{F}_{g}(V),\quad V\in\mathbb{R}^{T\times H\times W\times C}$ (1)
Meanwhile, for extracting video features, we use segment-wise Sampling following [31] to obtain fine-grained temporal cues. Given an input video $V\in\mathbb{R}^{T\times H\times W\times C}$, we divide it into $K$ segments, where each segment consists of $s=\frac{T}{K}$ frames. The video encoder $\mathcal{F}_{h}$, operates on low-resolution video segments $V_{k}\in\mathbb{R}^{s\times H\times W\times C}$ yielding global features that provide segment-wise temporal context.
$f_{h}=\mathcal{F}_{h}(V_{k}),\quad V_{k}\in\mathbb{R}^{s\times H\times W\times C}$ (2)
@@
$Z_{g}$ $=\mathcal{W}_{g}(f_{g}),\quad and\quad Z_{h}$ $=\mathcal{W}_{h}(f_{h})$ (3)
@@
Figure 3: Proposed Semi-automatic Annotation Pipeline. Our dataset for grounded conversation generation (GCG) is built from three video dataset types: i) Videos having masks only: Object patches are extracted from video frames using masks and processed by the Gemini-Pro model for initial object descriptions, which are then refined to produce detailed object captions. These refined captions and masks are again fed to Gemini-Pro model to create dense grounded captions. ii) Videos having bbox annotations and captions: Frames are first processed with a Video-LMM to generate a comprehensive caption which is combined with the original caption and fed to GPT-4o to obtain dense grounded captions. Masks are generated using frames and ground-truth bounding boxes with the SAM model. iii) Videos having object bboxes and referring expressions: Frames, bounding boxes, and referring expressions are input to GPT-4o for dense grounded captions, while masks are generated by feeding frames and bounding boxes to the SAM model.
@@ 3.3 Training Strategy
We train VideoGLaMM end-to-end in a single stage. As stated above, we use a dual encoder consisting of separate image and video encoders for processing spatial and temporal inputs to obtain local and global features, respectively. These encoders are initialized with weights of strong pre-trained encoders. During training, we keep the encoders fixed and only train the V→L adapters $\mathcal{W}_{g}$ and $\mathcal{W}_{h}$ associated with these encoders. These adapters are used to project the spatio-temporal visual features in the space of LLM and align the two modules. The spatio-temporal encoder is kept frozen and only the V→L adapters are updated. The textual features from the last layer of LLM, rich in spatial and temporal cues, are projected into the space of the pixel decoder using a multi-layer projection L→V adapter $\mathcal{W}_{p}$. For the LLM, we keep its weights frozen and only finetune LoRA [14] parameters during training. Both the frame encoder and pixel decoder are instantiated with pre-trained weights We keep the frame encoder and pixel decoder frozen and only train the L→V adapter layer. We optimize the output of the LLM by minimizing the cross entropy CE objective between the autoregressively obtained text output and dense grounded ground-truth caption. For the output of mask decoder, we optimize the intersection over union (IOU) between the predictions of mask decoder and ground-truth masks denoted as $\mathcal{L}_{masked}$. The total loss is the sum of CE loss and masked loss.
@@ 4 Our Benchmark & Annotation Pipeline
b) Videos with Bounding Box annotations and Captions: Fig. 3(b) shows the annotation process for the videos having both captions and object bounding box (Bbox) annotations. To obtain the corresponding dense grounded caption, the video frames are first passed to an open-source Video-LMM [27] to obtain a detailed caption, which is fed along with the reference ground truth caption to GPT-4o mini [34] to obtain the final dense grounded caption. The Bbox annotations are used as a prompts to SAM model [19] which takes the video frames as input and provides the masks corresponding to the objects.
@@ 5 Experimental Setup
Implementation details. Our spatio-temporal dual encoders follow the design of image and video encoders from [31]. For the image encoder, we use a pretrained CLIP ViT-L/14 (336 × 336)[37] model, and for the temporal encoder, we select the pretrained encoder of InternVideov2 (224 × 224) [50]. The V→L projectors are initialized with the weights of MLP adapter from [31]. The LLM is instantiated with Phi3-Mini-3.8B [1] weights. Both the frame encoder and pixel decoder are initialized with SAM2 [41] encoder-decoder weights. The training (Sec. 3.3) is carried out end-to-end on 4 Nvidia A100 40GB GPUs with a distributed training based on DeepSpeed [40].
Datasets. We train the model on our proposed grounded conversation (GCG) dataset containing 38k grounded video-QA triplets along with 83k objects and 671k fine-grained masks. During training, we also include a variety of other image and video segmentation datasets with our proposed benchmark dataset for more robust alignment. Our choice of image segmentation datasets include: ADE20K [55], COCO-Stuff [8], LVIS-PACO [38], refCOCO, refCOCO+, refCLEF, refCOCOg [15], LLaVA-Instruct-150k [25], ReasonSeg [20] and GranDf [39]. For video egmentation datasets we include train samples from Refer-DAVIS17 [17] and VideoInstruct100K [30].
Tasks. We evaluate VideoGLaMM on three challenging tasks: grounded conversation generation (GCG), visual grounding, and referring video segmentation. For grounded conversation generation, we curate a separate dataset of 308 refined video-QA triplets containing 826 objects and 22,762 fine-grained masks, following our proposed annotation pipeline. For Visual Grounding, we evaluate our model on challenging VidSTG [53] dataset, considering only the interrogative sentences as done by [32]. In the case of motion-guided video object segmentation, we leverage the MeViS [12] validation dataset. All the results on MeViS dataset are obtained via official CodaLab evaluation suite. We also report referring video object segmentation results on additional Ref-DAVIS-17 [18] dataset.
@@
5.1 Grounded Conversation Generation
The Grounded Conversation Generation (GCG) task aims to provide video-level detailed captions with specific phrases directly tied to corresponding segmentation masks in the video frames. For example, “<An adult> in white clothes holds a <cup> in the room", as shown in the first row of Fig. 4, features how each bracketed phrase is anchored to a unique segmentation mask. This creates a densely annotated caption that aligns textual descriptions with visual regions in the frames, enriching the video’s contextual interpretation. To obtain GCG output, we query the model with the following sample prompt: “Provide a detailed description of the image. Respond with interleaved segmentation masks for the corresponding parts of the answer.” The model generates a detailed caption along with interleaved segmentation masks, employing the format “<p>An adult woman in brown</p><SEG> is talking to another <p>adult man wearing jacket</p><SEG>” as shown in the third row of Fig. 4. We use special tokens, namely <p>, </p> and <SEG>, to delineate the start and end of each phrase and its corresponding region mask, respectively.
As shown in Table 1, our proposed Video-GLaMM performs better in generating detailed captions containing references to objects in the video frames, as is evident from high METEOR, CIDEr and CLAIR scores. Regarding the quality of masks, VideoGLaMM consistently outperforms baselines in terms of mIOU and Recall scores, signifying a higher overlap with ground-truth masks. Fig. 4 further shows the qualitative visualizations of VideoGLaMM on GCG samples.
@@ 5.2 Referring Video Segmentation
For referring video segmentation, the output should be grounded as per the given phrase, pointing towards specific instances in the video. Given a sentence or referring expression containing a specific object instance, the goal is to localize the object instances present across the video frames. This task operates in an open vocabulary setting, assessing the model’s ability to localize objects both spatially and temporally. Given a referring phrase expression Phrase, we prompt the model using the following instruction prompt to obtain the instance masks: “What is {Phrase} in this video? Respond with segmentation masks". Table 2 shows results on challenging MeViS dataset for motion-guided referring video segmentation. Both the region Jaccard $\mathcal{J}$ and boundary F-measure $\mathcal{F}$ are high in the case of VideoGLaMM, significantly outperforming the baselines. Similarly, the mean $\mathcal{J\&F}$ follows the same trend. Additionally, the scores corresponding to VideoLISA are reported with post-processing step. Notably, VideoLISA involves an additional post-processing step to boost performance. Therefore, we further fine-tune the VideoGLaMM on the task of referring segmentation post epoch 30 until epoch 40. Clearly, VideoGLaMM outperforms the Video-LISA (post-processed) on both $\mathcal{J}$ and $\mathcal{F}$, including the mean $\mathcal{J\&F}$. Additionally, VideoGLaMM outperforms baselines on Ref-DAVIS-17 dataset. The improved performance of VideoGLaMM can be credited to its training pipeline, which seamlessly integrates spatio-temporal dynamics into the model.
@@
Table 2: Performance comparison of VideoGLaMM on MeViS: VideoGLaMM shows superior performance on motion grounding and segmenting referring objects in the videos.
      Model       $\mathcal{J}$       $\mathcal{F}$       $\mathcal{J\&F}$
      LISA-7B [20]       58.4       54.9       61.9
      LISA-13B [20]       60.7       56.8       64.6
      TrackGPT-7B [58]       63.2       59.4       67.0
      TrackGPT-13B [58]       66.5       62.7       70.4
      VideoLISA [5]       68.8       64.9       72.7
      VideoGLaMM       69.5       65.6       73.3
Table 3: Performance comparison of VideoGLaMM on Ref-DAVIS-17: VideoGLaMM shows superior performance on segmenting referring objects in the videos.
@@ 5.3 Visual Grounding
To quantitatively assess VideoGLaMM’s visual grounding capability, we conduct quantitative evaluations on the benchmark test set of VidSTG dataset. The visual grounding task measures the adeptness of the model at correlating textual descriptions with visual elements in the video, a critical aspect of contextual comprehension. This ability is crucial for applications that integrate continuous visual data with language. The output of this task is refined masks that correlate with the given caption {caption}. To obtain the visual grounding output, we query the model with interrogative captions. For these captions, the prompt format follows “{caption} Please respond with a segmentation masks.". Table 4 shows VideoGLaMM’s improved visual grounding precision as it outperforms the baselines, demonstrating its fine-grained understanding of the videos.
@@
In addition to the above downstream tasks, in Sec. B of supplementary, we also integrate VideoGLaMM into a conditional video generation model [45]. VideoGLaMM provides temporally coherent masks that guides generative model in editing videos effectively. Please refer to Sections A and B of supplementary for additional quantitative and qualitative results respectively.
@@ 5.4 Ablation studies
Effect of Spatio-Temporal Dual Encoder. We employ separate image and video encoders to process spatial and temporal information. While spatial processing induces local information, temporal processing helps learn global features. Both are necessary from the perspective of grounding. To verify the effectiveness of dual spatio-temporal encoder, we conduct an ablation study where we measure the effectiveness of each encoder for grounded conversation generation (GCG) task (see Table 5). We notice that using only an image encoder gives suboptimal results, as we notice a drop in both the localization and captioning metrics. Using only video branch leads to the highest mIOU; however, relatively lower METEOR, CIDEr, and CLAIR scores. To obtain an optimal mIOU and good conversational abilities, VideoGLaMM uses both image and video encoder branches.
@@
Spatial vs Spatio-temporal Pixel decoder. Pixel decoder in VideoGLaMM can operate in two configurations. The first configuration processes video frames individually, ignoring temporal consistency. The second configuration employs both spatial and temporal branches for spatio-temporal context. Table 6 demonstrates the impact of spatiotemporal decoder on the GCG task. Results indicate that using only the spatial configuration reduces performance, with a nearly 3% drop in mIOU scores compared to the spatio-temporal configuration. Similarly, metrics like METEOR, CIDEr, and CLAIR also show a decline, underscoring the importance of using spatio-temporal configuration for pixel decoder.
@@
Effect of number of frames for Pixel Decoder. The pixel decoder receives the raw input frames encoded via frame encoder as input for predicting fine-grained grounded masks. During training, the pixel decoder also receives ground-truth masks which act as supervision signals. To provide more temporal supervision, we feed the pixel decoder with multiple input frames to enhance its temporal understanding. This allows it to learn semantic information that generalizes across frames. Table 7 shows the performance when 4 and 8 frames are input to the decoder. We observe that while the mIOU with 8 frames is slightly lower compared to 4 frames, the conversational quality measured by METEOR and CLAIR is higher. Hence, to achieve a decent mIOU with higher conversational output, we stick to 8 frames in the paper.
@@
Limitations and Future Work: Our GCG dataset plays a key role in enhancing the model’s grounding capabilities. While we validated annotations manually, some noise may still be present. Also, each scene contains several objects and the video descriptions do not exhaustively cover all objects in the scenes. A higher-quality densely annotated set could further boost model performance but would require substantial annotation resources. Additionally, VideoGLaMM struggles with objects of varying granularities, likely due to limited representation in the training data. Another improvement is to extend VideoGLaMM for longer videos, as the current GCG dataset mainly focuses on short-medium duration clips.